# Gradient-free training of recurrent neural networks

### v1

@@ H.1.2 Hyperparameters
Van der Pol 1D Van der Pol Lorenz Rössler forced Van der Pol Weather
@@ H.2 ESN
For the Van der Pol problem many hyperparameter combinations were tried out until we were able to find a model with good performance and sufficient robustness to a change in the random seed. The hyperparameter combinations we considered are shown in Table 4. Due to the high expenses of a grid-search approach, a random search on a grid was employed and 1000 hyperparameter combinations were considered. The hyperparameter choice with the best performance on the validation dataset was selected and then evaluated on the test dataset. The final choice of hyperparameters is documented in Table 3, and any unmentioned hyperparameters are assumed to be set to the default value from the reservoirpy library.
@@ H.3.1 Hyperparameters
We used the clipped shPLRNN trained by GTF. For the Van der Pol, Lorenz and the weather datasets we considered a fixed GTF parameter $\alpha$, while for the Rössler we considered an adaptive $\alpha$ (starting from an upper bound) as proposed by Hess et al. [2023]. The code repository by Hess et al. [2023] was used to perform the computations. The hyperparameters selected for all datasets are detailed in Table 5. Any hyperparameters not specified are set to their default values in the corresponding repository of Hess et al. [2023]. For Rössler dataset, finding optimal hyperparameters was more challenging, and for training, we also utilized regularizations for the latent and observation models.

### v2

@@ H.1.2 Hyperparameters
Van der Pol 1D Van der Pol Lorenz Rössler force Van der Pol Weather
@@ H.2 ESN
For the Van der Pol problem many hyperparameter combinations were tried out until we were able to find a model with good performance and sufficient robustness to a change in the random seed. The hyperparameter combinations we considered are shown in Table 4. Due to the high expenses of a grid-search approach, a random search was employed and 1000 hyperparameter combinations were considered. The hyperparameter choice with the best performance on the validation dataset was selected and then evaluated on the test dataset. The final choice of hyperparameters is documented in Table 3, and any unmentioned hyperparameters are assumed to be set to the default value from the reservoirpy library.
@@ H.3.1 Hyperparameters
We used the clipped shPLRNN trained by GTF. For the Van der Pol, Lorenz and the weather datasets we considered a fixed GTF parameter $\alpha$, while for the Rössler we considered an adaptive $\alpha$ (starting from an upper bound) as proposed by [Hess et al., 2023]. The code repository by [Hess et al., 2023] was used to perform the computations. The hyperparameters selected for all datasets are detailed in Table 5. Any hyperparameters not specified are set to their default values in the corresponding repository of [Hess et al., 2023]. For Rössler dataset, finding optimal hyperparameters was more challenging, and for training, we also utilized regularizations for the latent and observation models.
@@ H.4.2 Hardware
Appendix I Additional experiment with real-world data
We provide two additional experiments to showcase our the performance of our model on real world data. From our experience it was not necessary to perform exhaustive hyperparamter tuning in order to obtain decent performance from a sampled RNN model. The same hardware was used as for the weather data experiment mentioned in Section H.4.2.
Electricity consumption
We use the individual power consumption dataset by Hebrail and Berard [2006] and predict the voltage feature. This dataset contains measurements made in a one-minute interval, and we consider a period of four weeks as our dataset. Two weeks are used as training data, one week as validation and one as test data, as shown in Figure 13. We added sin and cos time-embeddings of hour and day. We use 240 time-delay embeddings (amounting to 4 hours) and always predict a horizon of 120 steps (2 hours). The sampled RNN has 128 hidden nodes and a $tanh$ activation. The results are plotted in Figure 14. Training time was 6.5 seconds. The MSE error on training, validation and test data is 1.565, 1.427 and 1.428 V respectively.
Figure 13: Individual household electrical power consumption data split into training, validation and testing portions.
Figure 14: Test set predictions of individual household electricity consumption.
Bike traffic
We use a dataset containing bike traffic counts in Copenhagen bik . This dataset contains only the bike count and the date and time information. We use the bike count as a target variable. We added sin and cos time-embeddings of hour, day, and month. The largest portion of data without missing values is used as the dataset, and split into training, validation and test as shown in Figure 15. We use a sampled RNN with 256 hidden nodes and $tanh$ activation. A time delay of 14 hours is used, and the prediction horizon is 7 hours. Training time was 0.65 seconds. The MSE error on the training, validation and test set is 90, 160, 140, respectively.
Figure 15: Bike count data split into training, validation and testing portions.
Figure 16: Test set predictions of bike counts.
Appendix J Comparison without SWIM and without Koopman
In this section, we report on some results performed to separate each contribution of the parts we use in this paper, namely the SWIM sampling strategy and the Koopman operator. We ran two settings (excluding our method already reported). One is with the addition of the Koopman operator and EDMD, but where we sample the weights of the hidden layer with a standard Gaussian distribution and the biases with a uniform distribution. This sampling strategy is commonly used for random neural networks, which is why we chose these distributions in particular. In the second setting, we exclude the EDMD step and simply project down the original state space after applying the hidden layer. The results can be found in Table 8. We observe that the model that incorporates both Koopman and SWIM performs best over all the experiments we tested, and with typically smaller error bars.
Table 8: Test setting (A): without SWIM and with Koopman, (B): with SWIM and without Koopman, (C) with SWIM and with Koopman (our). Results for Van der Pol are reported with MSE and the chaotic systems Lorenz and Rössler are reported with EKL, and all experiments report mean, minimum, and maximum respectively.
Example (A) (B) (C)
Van der Pol 2.90e-3 (6e-4,8.1e-3) 3.12e-2 (3.05e-2, 3.16e-2) 9.55e-4 (7.08e-4, 1.28e-3)
Lorenz 7.79e-3 (5.79e-3, 9.87e-3) 8.41e-3 (6.45e-3, 1.02e-2) 4.36e-3 (3.66e-3, 5.36e-3)
Rössler 4.47e-1 (1.7e-4, 2.21e-0) 7.51e-3 (1.00e-4,3.63e-2) 1.57e-4 (5.86e-5, 3.82e-4)
Finally, to see how our method compares to one without Koopman when one extends the size of the time step to predict the Van der Pol system, we ran an experiment for time steps $\Delta t=[0.1,0.2,0.3,0.4,0.5]$, and the results can be found in Figure 17. Here, we observe that the method, including Koopman, consistently performs better and has a much more stable error bar as we increase $\Delta t$. This has also been observed separately (and without our special dictionary / RNN setting) by Mauroy and Goncalves [2017].
(a) without SWIM
(b) using SWIM
Figure 17: Running our model and a model without Koopman while increasing the size of the time step the models predict, $\Delta t=[0.1,0.2,0.3,0.4,0.5]$, results reported on test set with MSE and min/max error bars.